# Exploring the Strongly-Interacting Regime of Effective Multi-Body Interactions in a Trapped Ultracold Atom System

### v2

@@ Abstract
A two-body interaction or force between quantum particles is ubiquitous in nature, and the microscopic description in terms of the bare two-body interaction is the basis for quantitatively describing interacting few- and many-body systems. Alternatively, the effective description in terms of an effective two-body interaction successfully captures the essence of the systems. However, for several important observations, the explanation in terms of an effective two-body interaction is not satisfactory, and the effective three-body interaction has played an essential role in understanding the systems. In this study, we investigate a few-body system comprising of ultracold bosons tightly confined in a deep optical lattice site, which is effectively described as zero-dimensional bosons. By combining an occupancy-resolving high-resolution laser spectroscopy with an inter-orbital Feshbach resonance controlling the bare two-body interaction over a wide range, we experimentally reveal the behaviors of few-atom systems in a strongly-interacting regime. Our results, for which perturbative calculations do not provide proper explanations, serve as a valuable and precise benchmark for theoretical approaches to strongly interacting few-body systems. As one important illustration, we obtain a clear signature of an effective four-body interaction, evidenced by the binding energies of four and more atoms. This work is an important first step for our deeper understanding of an important relation to a strongly-interacting mesoscopic Bose polaron problem.
@@ I Introduction
A few-body quantum system is usually defined as the one which consists of a small number of well-defined constituent particles, often trapped or self-trapped under some confining potential [1]. Few-quark states, light nuclear and hadronic systems, few-electron atomic and molecular systems, quantum dots, and trapped cold-atom systems are the typical examples and have recently been studied extensively and intensively [2, 3, 4]. In these few-body systems, a two-body interaction or force between constituent quantum particles, usually described by scattering lengths at a low collision energy, plays a dominant role in the quantitative description of the system.
@@
Recent progress of ultracold atoms has also provided a useful platform for studying few-body atomic systems. High controllability and a rich variety of powerful spectroscopic techniques of this cold-atom system enables the investigation of, for example, a novel phenomenon of the Efimov effect in which an effective three-body interaction emerges via multiple scatterings via resonant two-body interactions, resulting in a series of universal weakly-bound states characterized by a three-body parameter. Similarly, four- and more atom bound states are observed and explained by using a three-body parameter or potential [18, 19]. Another important example of few-body atomic system is a well-localized few atom system tightly confined in each of an optical lattice potential, and the effective three-body interaction were extensively studied with various methods such as quantum phase collapse and revival technique [20], lattice modulation spectroscopy [21], high-resolution clock spectroscopy both for bosons [22, 23, 24] and fermions [25], and the results are quantitatively explained by perturbative calculations [26, 27, 28].
The procedure of the determination of the energies with the perturbation theory is elaborated on in Ref. [27]. The emergence of the effective interactions is understood in the physical picture based on the microscopic description by integrating out the degrees of freedom of virtual excitations to higher vibrational states. Here, we can find an analogy with the effective three-body force of Fujita-Miyazawa type in nuclear physics, where nucleons, a bare two-body nuclear force, and the $\Delta$ excitations correspond to the atoms in the vibrational ground state of a harmonic trap, the bare two-body interatomic force, and the vibrational excitations of the atoms, respectively.
@@
Effective multi-body interactions have also been of considerable interest in different contexts of the study of localized spin models. Effective multi-spin interactions, arising from higher-order processes of pairwise exchange interactions, was theoretically proposed to explain the magnetic behaviors in solid helium [33] and some organic compounds [34]. In addition, similar effective multi-spin interactions have been recently engineered in artificial systems of a chain of ions in an ion trap [35] and localized spins loaded into an optical lattice with unit filling [36], which would play an important role in quantum computation and quantum information. Moreover, an effective three-body interaction for photon is successfully implemented by exploiting Rydberg interactions [37].
@@ III Effective multi-body interactions with perturbation theory
Before the discussion of the experiment, we here present the perturbation theory of the effective multi-body interactions studied in this work, the results of which are later compared with the experimental observations to illustrate the novel features of the strongly interacting regime of the multi-body interactions. The system consisting of multiple 174Yb atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}$ states is described by Eq. (II). The trap angular frequency $\omega$ is set to be $\omega_{g}=2\pi\times 35.9$ kHz if all $N$ atoms are in the ground state, but it is assumed to be an average value of
@@
$\quad-6\alpha_{4,3}^{(3)}\left(\frac{a_{gg}}{l_{g}}\right)^{3}\hbar\omega_{g}-18\alpha_{5}^{(3)}\left(\frac{a_{gg}}{l_{g}}\right)^{3}+O(a^{4})$ (6)
@@
By computing the left-hand side up to the third order in scattering lengths, the effective three- and four-body interaction energies are found and their explicit formulae are provided by Eqs. (A) and (A) of Appendix A.
@@
for $U_{gg}=2\pi\hbar\times 4.04$ kHz, whereas the explicit formulae of $U_{egg}$ and $U_{eggg}$ are provided by Eqs. (26) and (27) of Appendix A. The resulting $\nu_{3}-\nu_{1}=[2(U_{eg}-U_{gg})+U_{egg}-U_{ggg}]/(2\pi\hbar)$ as a function of $\nu_{2}-\nu_{1}=(U_{eg}-U_{gg})/(2\pi\hbar)$ and $\nu_{4}-\nu_{1}=[3(U_{eg}-U_{gg}+U_{egg}-U_{ggg})+U_{eggg}-U_{gggg}]/(2\pi\hbar)$ as a function of $\nu_{3}-\nu_{2}=(U_{eg}-U_{gg}+U_{egg}-U_{ggg})/(2\pi\hbar)$ under the variation of $U_{eg}$ in the equal frequency approximation are plotted in Figs. 4(a) and 5(a), respectively, which are compared with the experimental data later in Sec. V.
@@ V Experimental Results
Figure 3: Variation of resonance positions for $n$-occupancy sites via a Feshbach resonance. a, Dependence on a magnetic field. The resonance positions corresponding to the $n=1$ (gray square), $n=2$ (blue circle), $n=3$ (red triangle), $n=4$ (green diamond), and $n=5$ (orange inverted triangle) sites obtained from the spectra at each magnetic field strength are plotted. Note that the origin of the vertical frequency axis is taken as the resonance frequency of the single occupancy (horizontal dashed line). The error bars show the standard deviation of the resonance positions obtained from at least three independent scans, including errors associated with the correction for long-term laser frequency drift (see Appendix C). The vertical dotted line denotes the Feshbach resonance point between ${{}^{1}}S_{0}\leftrightarrow{{}^{3}}P_{2}\ (m_{J}=+2)$ atoms at a magnetic field strength of 398 mG, obtained from the measured $n=2$ resonance positions (see Appendix D). Note that calibration of the magnetic field is done by observing the resonance position of ${{}^{1}}S_{0}\leftrightarrow{{}^{3}}P_{2}\ (m_{J}=\pm 2)$, and the accuracy of the magnetic field is about few mG. The solid curves show calculated resonance positions including only the two-body interaction terms, where the solid curve at $n=2$ (blue) shows a result of the theoretical calculation and the solid curves at $n=3,4$ and $5$ (red, green and orange) shows a result of the solid curve at $n=2$ (blue) multiplied by 2, 3 and 4, respectively (see Appendices D and E). b, Dependence on the scattering length $a_{eg}$. The legends are common to those of a. The horizontal axis represents the scattering length $a_{eg}$ for atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}\ (m_{J}=+2)$ states, divided by the harmonic oscillator length $l_{eg}=39$ nm corresponding to the optical lattice depth of $25\,E_{R}$. The blue solid curve at the $n=2$ plots shows a result of the theoretical calculation. The horizontal error bars are evaluated for each field strength from the error in the determination of the scattering length (see Appendix D).
@@
In order to enter the strongly-interacting regime beyond the perturbative treatment of the interatomic interaction, we combine the inter-orbital Feshbach resonance with the occupancy-resolving high-resolution spectroscopy. Here, the scattering length $a_{eg}$ between the 174Yb atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}~{}(m_{J}=+2)$ states is controlled with a magnetic field via an anisotropy-induced Feshbach resonance [23]. Figure 3 shows the measured frequency shift of the resonance position of the multiply occupied site with respect to the resonance frequency of the singly occupied site, as a function of (a) a magnetic field and (b) the scattering length which is determined from the $n=2$ data (see Appendix D for the details of the calculation). As the magnetic field approaches the value at the Feshbach resonance point indicated by the dotted line (see Appendix D), the frequency shift also increases, reflecting the larger absolute values of the scattering length $a_{eg}$. We also note that the resonance signals corresponding to the multiple occupancies, $n=5$ in particular, become weaker around the Feshbach resonance, possibly due to the spectral broadening caused by the fast three-body inelastic loss of the atoms in the excited state.
@@
Since the strong interactions are induced via the inter-orbital Feshbach resonance between $(n-1)$ |$\ket{g}$-state bosons and a single |$\ket{e}$-state atom, rather than among $n$ |$\ket{g}$-state bosons, the injected single |$\ket{e}$-state atom introduces the modification of the atomic wavefunctions of the relative motion between the |$\ket{e}$-state and other |$\ket{g}$-state atoms, resulting in the non-linearity of interactions. In addition, it is instructive to consider this injected few-body state in the context of the Bose polaron problem [40, 41] by regarding a single |$\ket{e}$-state atom as an impurity, which may provide another indispensable insight for this system.
@@
An interesting feature is found in the comparison of the experimental data with the perturbative calculations described in Sec. III. The dashdot and solid curves in Fig. 4(a) show the results of perturbative calculations up to the second and third orders of the two-body interaction energy, respectively (see Eqs. (14) and (26)). Clearly, the second-order perturbation calculation does not reproduce any of the experimental data observed in our experiment on the one hand. The third-order result, on the other hand, can properly reproduce the data corresponding to the small scattering lengths where the perturbative calculation is justified. Note that the experimental data are all located on the upper side of the line of Eq. (16), which indicates the inequality $U_{egg}>U_{ggg}$ for the effective three-body interactions and the amplitude of the effective three-body interactions $U_{egg}-U_{ggg}$ is plotted in Fig. 4(b) as a function of $a_{eg}/l_{eg}$. This feature is properly captured by the perturbation calculation up to the third order, although the magnitudes of the deviations are not.
@@
Figure 4: Effective three-body interaction in strongly-interacting regime. a, Evidence of an effective three-body interaction in strongly-interacting regime obtained for the $n=3$ resonance. The horizontal and vertical axes show the frequency shifts of $n=2$ and $n=3$ with respect to the resonance position of $n=1$, respectively. The colorbar shows the scattering length $a_{eg}$ divided by the harmonic oscillator length $l_{eg}$. The error bars for each axis are the same as those in Fig. 3. The dashed line represents a linear relation with slope 2, which holds between $\nu_{3}-\nu_{1}$ and $\nu_{2}-\nu_{1}$ when only the effective two-body interactions are incorporated, and the deviation from this line in the plots directly indicates the presence of effective three-body interactions. The dashdot and solid curves show the results of the perturbative calculations up to the second and third orders of the two-body interaction energy, respectively. b, Amplitude of effective three-body interactions. Plotted here are values of $(U_{egg}-U_{ggg})/h=(\nu_{3}-\nu_{1})-2(\nu_{2}-\nu_{1})$ (Eq. (63)) from measured resonance positions of $n=2$ and $3$. The error bars are calculated with the error propagation formula by using the error bars in Fig. 4(a).
@@
Figure 5: Evidence of an effective four-body interaction. a, Evidence of an effective four-body interaction obtained for the $n=4$ resonance. The horizontal and vertical axes represent the difference between the $n=2$ and $n=3$ resonance positions and that between $n=1$ and $n=4$, respectively. The colorbar shows the scattering length $a_{eg}$ divided by the harmonic oscillator length $l_{eg}$. Error bars for the vertical axis are the same as those in Fig. 3, and error bars for the horizontal axis represent the standard deviation obtained in the analysis with respect to the resonance position of $n=2$ (see Appendix C). The dashed line represents a linear relation with slope 3, which holds between $\nu_{4}-\nu_{1}$ and $\nu_{3}-\nu_{2}$ when only up to effective three-body interactions are incorporated, and the deviation from this line in the plots is the direct evidence of the presence of effective four-body interactions. The solid curve shows the result of the perturbative calculation up to the third order of the two-body interaction energy. b, Amplitude of effective four-body interactions. Plotted here are values of $(U_{eggg}-U_{gggg})/h=(\nu_{4}-\nu_{1})-3(\nu_{3}-\nu_{2})$ (Eq. (64)) from measured resonance positions of $n=2$, $3$, and $4$. The error bars are calculated with the error propagation formula by using the error bars in Fig. 5(a). c, Contribution of effective four-body interactions to the $n=5$ resonance. Plotted here are the measured $n=5$ resonance positions (orange inverted triangles), the calculated $n=5$ resonance positions by incorporating only up to effective three-body interactions (3-BI, red circles), and the calculated $n=5$ resonance positions by incorporating up to effective four-body interactions (4-BI, green squares). Note that the plots at each scattering length are slightly shifted horizontally to avoid the overlap of the symbols. The error bars for the measured $n=5$ resonance positions are the same as the error bars in Fig. 3, and the error bars in the calculated values are obtained with the error propagation formula by using the error bars for $n=2,3$, and $4$ resonance positions.
@@
Note that the plots are all located on the lower side of the line, which shows that the inequality $U_{eggg}<U_{gggg}$ holds for the effective four-body interactions, where $U_{gggg}$ and $U_{eggg}$ represent the interaction energy for the four atoms in the ground state |$\ket{g}$ and that for three in the ground state and one in the excited state |$\ket{e}$, respectively, and the amplitude of the effective four-body interactions $U_{eggg}-U_{gggg}$ is plotted in Fig. 5(b) as a function of $a_{eg}/l_{eg}$. Since the Feshbach resonance is only induced between the atoms in the ground and excited states and not between the ground state atoms, $U_{eggg}$ should be responsible for the emergence of the effective four-body interaction.
Figure 5(a) also shows as a solid curve the result of the perturbative calculation up to the third order of the two-body interaction energy, valid only in a region of relatively small scattering length (see Eqs. (15) and (27)). Clearly, the perturbative calculations do not reproduce the experimental data including the direction of the deviation from the linear relationship Eq. (18), indicating the non-perturbative nature of the observed effective four-body interaction.

### v4

@@ Abstract
A two-body interaction or force between quantum particles is ubiquitous in nature, and the microscopic description in terms of the bare two-body interaction is the basis for quantitatively describing interacting few- and many-body systems. Alternatively, the effective description in terms of an effective two-body interaction successfully captures the essence of the systems. However, for several important observations, the explanation in terms of an effective two-body interaction is not satisfactory, and the effective three-body interaction has played an essential role in understanding the systems. In this study, we investigate a few-body system comprising of ultracold bosons tightly confined in a deep optical lattice site, which is effectively described as zero-dimensional bosons. By combining an occupancy-resolving high-resolution laser spectroscopy with an inter-orbital Feshbach resonance controlling the bare two-body interaction over a wide range, we experimentally reveal the behaviors of few-atom systems in a strongly-interacting regime. Our results, for which perturbative calculations do not provide proper explanations, serve as a valuable and precise benchmark for theoretical approaches to strongly-interacting few-body systems. As one important illustration, we obtain a clear signature of an effective four-body interaction, evidenced by the binding energies of four and more atoms. This work is an important step for our deeper understanding of strongly-interacting few-body systems.
@@ I Introduction
A few-body quantum system is usually defined as the one which consists of a small number of well-defined constituent particles, often trapped or self-trapped under some confining potential [1]. Few-quark states, light nuclear and hadronic systems, few-electron atomic and molecular systems, quantum dots, and trapped cold-atom systems are the typical examples, which together with universal aspects across them have recently been studied intensively [2, 3, 4, 5, 6, 7, 8]. In these few-body systems, a two-body interaction or force between constituent quantum particles, usually described by scattering lengths at a low collision energy, plays a dominant role in the quantitative description of the system.
@@
Recent progress of ultracold atoms has also provided a useful platform for studying few-body atomic systems. High controllability and a rich variety of powerful spectroscopic techniques of this cold-atom system enables the investigation of, for example, a novel phenomenon of the Efimov effect in which an effective three-body interaction emerges via resonant two-body interactions, resulting in a series of universal weakly-bound states characterized by a three-body parameter [22, 23, 24, 25]. Similarly, four- and more atom bound states are observed and explained by using a three-body parameter or potential [26, 27]. Another important example of few-body atomic system is a well-localized few atom system tightly confined in each of an optical lattice potential, and the effective three-body interaction were extensively studied with various methods such as quantum phase collapse and revival technique [28], lattice modulation spectroscopy [29], and high-resolution clock spectroscopy both for bosons [30, 31, 32] and fermions [33]. Such progress on the experimental side has stimulated further development on the theoretical side, including exact solutions to two atoms in a harmonic trap [34, 35, 36, 37, 38], perturbative calculations of effective multi-body interactions [39, 40, 41], and analytical and numerical approaches to strongly-interacting few-body systems [2, 3, 4, 5, 6, 7, 8].
The emergence of the effective multi-body interactions is understood in the physical picture based on the microscopic description by integrating out the degrees of freedom of virtual excitations to higher vibrational states. Here, we can find an analogy with the effective three-body force of Fujita-Miyazawa type in nuclear physics, where nucleons, a bare two-body nuclear force, and the $\Delta$ excitations correspond to the atoms in the vibrational ground state of a harmonic trap, the bare two-body interatomic force, and the vibrational excitations of the atoms, respectively.
@@
Effective multi-body interactions have also been of considerable interest in different contexts of the study of localized spin models. Effective multi-spin interactions, arising from higher-order processes of pairwise exchange interactions, were theoretically proposed to explain the magnetic behaviors in solid helium [45] and some organic compounds [46]. In addition, similar effective multi-spin interactions have been recently engineered in artificial systems of a chain of ions in an ion trap [47] and localized spins loaded into an optical lattice with unit filling [48], which would play an important role in quantum computation and quantum information. Moreover, an effective three-body interaction for photon is successfully implemented by exploiting Rydberg interactions [49].
@@ III Effective multi-body interactions with perturbation theory
Before the discussion of the experiment, we here present the perturbation theory of the effective multi-body interactions studied in this work, the results of which are later compared with the experimental observations to illustrate the novel features of the strongly-interacting regime of the multi-body interactions. The system consisting of multiple 174Yb atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}$ states is described by Eq. (II). The trap angular frequency $\omega$ is set to be $\omega_{g}=2\pi\times 35.9$ kHz if all $N$ atoms are in the ground state, but it is assumed to be an average value of
@@
$\quad-6\alpha_{4,3}^{(3)}\left(\frac{a_{gg}}{l_{g}}\right)^{3}-18\alpha_{5}^{(3)}\left(\frac{a_{gg}}{l_{g}}\right)^{3}+O(a^{4})$ (6)
@@
By computing the left-hand side up to the third order in scattering lengths, the effective three- and four-body interaction energies are found and their explicit formulae are provided by Eqs. (A) and (A) of Appendix A. We note that the effective range does not appear up to the order we are working on [40].
@@
for $U_{gg}=2\pi\hbar\times 4.04$ kHz, whereas the explicit formulae of $U_{egg}$ and $U_{eggg}$ are provided by Eqs. (26) and (27) of Appendix A. The resulting $\nu_{3}-\nu_{1}=[2(U_{eg}-U_{gg})+U_{egg}-U_{ggg}]/(2\pi\hbar)$ as a function of $\nu_{2}-\nu_{1}=(U_{eg}-U_{gg})/(2\pi\hbar)$ and $\nu_{4}-\nu_{1}=[3(U_{eg}-U_{gg}+U_{egg}-U_{ggg})+U_{eggg}-U_{gggg}]/(2\pi\hbar)$ as a function of $\nu_{3}-\nu_{2}=(U_{eg}-U_{gg}+U_{egg}-U_{ggg})/(2\pi\hbar)$ under the variation of $U_{eg}$ in the equal frequency approximation are plotted in Figs. 4(a,b) and 5(a,b), respectively, which are compared with the experimental data later in Sec. V.
@@ V Experimental Results
Figure 3: Variation of resonance positions for $n$-occupancy sites via a Feshbach resonance. a, Dependence on a magnetic field. The resonance positions corresponding to the $n=1$ (gray square), $n=2$ (blue circle), $n=3$ (red triangle), $n=4$ (green diamond), and $n=5$ (orange inverted triangle) sites obtained from the spectra at each magnetic field strength are plotted. Note that the origin of the vertical frequency axis is taken as the resonance frequency of the single occupancy (horizontal dashed line). The error bars show the standard deviation of the resonance positions obtained from at least three independent scans, including errors associated with the correction for long-term laser frequency drift (see Appendix C). The vertical dotted line denotes the Feshbach resonance point between ${{}^{1}}S_{0}\leftrightarrow{{}^{3}}P_{2}\ (m_{J}=+2)$ atoms at a magnetic field strength of 398 mG, obtained from the measured $n=2$ resonance positions (see Appendix D). Note that calibration of the magnetic field is done by observing the resonance position of ${{}^{1}}S_{0}\leftrightarrow{{}^{3}}P_{2}\ (m_{J}=\pm 2)$, and the accuracy of the magnetic field is about few mG. The solid curves show calculated resonance positions including only the two-body interaction terms, where the solid curve at $n=2$ (blue) shows a result of the theoretical calculation. The solid curves at $n=3,4$ and $5$ (red, green and orange) show a result of the solid curve at $n=2$ (blue) multiplied by 2, 3 and 4, respectively, which for $n$ occupancies are explicitly given by $(n-1)(U_{eg}-U_{gg})$ (see Appendices D and E). We note that the relatively large deviation of the measured resonance positions from the solid curves, even in a regime where $a_{eg}$ is relatively small, is due to neglecting the contributions of effective three- and four-body forces. Note that the experimental data in the perturbative regime in Fig. 4(a) and Fig. 5(a) are close to the calculation, where the effective three- and four-body forces are perturbatively taken into consideration. b, Dependence on the scattering length $a_{eg}$. The legends are common to those of a. The horizontal axis represents the scattering length $a_{eg}$ for atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}\ (m_{J}=+2)$ states, divided by the harmonic oscillator length $l_{eg}=39$ nm corresponding to the optical lattice depth of $25\,E_{R}$. The blue solid curve at the $n=2$ plots shows a result of the theoretical calculation. The horizontal error bars are evaluated for each field strength from the error in the determination of the scattering length (see Appendix D). Note that the energy shift of about $-3$ kHz for $a_{eg}=0$ in b is due to the non-zero value of the interaction energy between the ground-state atoms ($U_{gg}$).
@@
In order to enter the strongly-interacting regime beyond the perturbative treatment of the interatomic interaction, we combine the inter-orbital Feshbach resonance with the occupancy-resolving high-resolution spectroscopy. Here, the scattering length $a_{eg}$ between the 174Yb atoms in the ground ${{}^{1}}S_{0}$ and excited ${{}^{3}}P_{2}~{}(m_{J}=+2)$ states is controlled with a magnetic field via an anisotropy-induced Feshbach resonance [31]. Figure 3 shows the measured frequency shift of the resonance position of the multiply occupied site with respect to the resonance frequency of the singly occupied site, as a function of (a) a magnetic field and (b) the scattering length which is determined from the $n=2$ data (see Appendix D for the details of the calculation). As the magnetic field approaches the value at the Feshbach resonance point indicated by the dotted line (see Appendix D), the frequency shift also increases, reflecting the larger absolute values of the scattering length $a_{eg}$. We note that the resonance signals corresponding to the multiple occupancies, $n=5$ in particular, become weaker around the Feshbach resonance, possibly due to the spectral broadening caused by the fast three-body inelastic loss of the atoms in the excited state. We also note that the three-dimensional lattice is considered as an ideally isolated matrix of zero-dimensional systems even in the strongly interacting regime, since the band widths for the ground as well as higher bands are about one-order-of-magnitude smaller than the band gaps and the interatomic interaction energy.
@@
Since the strong interactions are induced via the inter-orbital Feshbach resonance between $(n-1)$ |$\ket{g}$-state bosons and a single |$\ket{e}$-state atom, rather than among $n$ |$\ket{g}$-state bosons, the injected single |$\ket{e}$-state atom introduces the modification of the atomic wavefunctions of the relative motion between the |$\ket{e}$-state and other |$\ket{g}$-state atoms, resulting in the non-linearity of interactions. In addition, it is instructive to consider this injected few-body state in the context of the Bose polaron problem [52, 53] by regarding a single |$\ket{e}$-state atom as an impurity.
@@
An interesting feature is found in the comparison of the experimental data with the perturbative calculations described in Sec. III. The dashdot and solid curves in Fig. 4(a) show the results of perturbative calculations up to the second and third orders of the two-body interaction energy, respectively (see Eqs. (14) and (26)). Clearly, the second-order perturbation calculation does not reproduce any of the experimental data observed in our experiment on the one hand. The third-order result, on the other hand, can properly reproduce the data corresponding to the small scattering lengths where the perturbative calculation is justified. Note that the experimental data are all located on the upper side of the line of Eq. (16), which indicates the inequality $U_{egg}>U_{ggg}$ for the effective three-body interactions and the amplitude of the effective three-body interactions $U_{egg}-U_{ggg}$ is plotted in Fig. 4(b) as a function of $U_{eg}/\hbar\omega_{eg}$. This feature is properly captured by the perturbation calculation up to the third order, although the magnitudes of the deviations are not.
@@
Figure 4: Effective three-body interaction in strongly-interacting regime. a, Evidence of an effective three-body interaction in strongly-interacting regime obtained for the $n=3$ resonance. The horizontal and vertical axes show the frequency shifts of $n=2$ and $n=3$ with respect to the resonance position of $n=1$, respectively. The colorbar shows the two-body interaction energy $U_{eg}$ divided by $\hbar\omega_{eg}$. The error bars for each axis are the same as those in Fig. 3. The dashed line represents a linear relation with slope 2, which holds between $\nu_{3}-\nu_{1}$ and $\nu_{2}-\nu_{1}$ when only the effective two-body interactions are incorporated, and the deviation from this line in the plots directly indicates the presence of effective three-body interactions. b, Amplitude of effective three-body interactions. Plotted here are values of $(U_{egg}-U_{ggg})/h=(\nu_{3}-\nu_{1})-2(\nu_{2}-\nu_{1})$ (Eq. (63)) from measured resonance positions of $n=2$ and $3$. The error bars are calculated with the error propagation formula by using the error bars in Fig. 4(a). In both a and b, the dashdot and solid curves show the results of the perturbative calculations up to the second and third orders of the two-body interaction energy, respectively.
@@
Figure 5: Evidence of an effective four-body interaction. a, Evidence of an effective four-body interaction obtained for the $n=4$ resonance. The horizontal and vertical axes represent the difference between the $n=2$ and $n=3$ resonance positions and that between $n=1$ and $n=4$, respectively. The colorbar shows the two-body interaction energy $U_{eg}$ divided by $\hbar\omega_{eg}$. Error bars for the vertical axis are the same as those in Fig. 3, and error bars for the horizontal axis represent the standard deviation obtained in the analysis with respect to the resonance position of $n=2$ (see Appendix C). The dashed line represents a linear relation with slope 3, which holds between $\nu_{4}-\nu_{1}$ and $\nu_{3}-\nu_{2}$ when only up to effective three-body interactions are incorporated, and the deviation from this line in the plots is the direct evidence of the presence of effective four-body interactions. b, Amplitude of effective four-body interactions. Plotted here are values of $(U_{eggg}-U_{gggg})/h=(\nu_{4}-\nu_{1})-3(\nu_{3}-\nu_{2})$ (Eq. (64)) from measured resonance positions of $n=2$, $3$, and $4$. The error bars are calculated with the error propagation formula by using the error bars in Fig. 5(a). In both a and b, the solid curve shows the result of the perturbative calculation up to the third order of the two-body interaction energy. c, Contribution of effective four-body interactions to the $n=5$ resonance. Plotted here are the measured $n=5$ resonance positions (orange inverted triangles), the calculated $n=5$ resonance positions by incorporating only up to effective three-body interactions (3-BI, red circles), and the calculated $n=5$ resonance positions by incorporating up to effective four-body interactions (4-BI, green squares). Note that the plots at each scattering length are slightly shifted horizontally to avoid the overlap of the symbols. The error bars for the measured $n=5$ resonance positions are the same as the error bars in Fig. 3, and the error bars in the calculated values are obtained with the error propagation formula by using the error bars for $n=2,3$, and $4$ resonance positions.
@@
Note that the plots are all located on the lower side of the line, which shows that the inequality $U_{eggg}<U_{gggg}$ holds for the effective four-body interactions, where $U_{gggg}$ and $U_{eggg}$ represent the interaction energy for the four atoms in the ground state |$\ket{g}$ and that for three in the ground state and one in the excited state |$\ket{e}$, respectively, and the amplitude of the effective four-body interactions $U_{eggg}-U_{gggg}$ is plotted in Fig. 5(b) as a function of $U_{eg}/\hbar\omega_{eg}$. Since the Feshbach resonance is only induced between the atoms in the ground and excited states and not between the ground state atoms, $U_{eggg}$ should be responsible for the variation of the effective four-body interaction.
Figure 5(a,b) also show as a solid curve the result of the perturbative calculation up to the third order of the two-body interaction energy, valid only in a region of relatively small scattering length (see Eqs. (15) and (27)). Clearly, the perturbative calculations do not reproduce the experimental data including the direction of the deviation from the linear relationship Eq. (18), indicating the non-perturbative nature of the observed effective four-body interaction. Note that the possible effects coming from non-harmonic potential terms, as well as the uncertainties due to the magnetic field calibration, $a_{gg}$, the lattice depth calibration, and the trap frequencies, are negligible in our conditions. These error budgets are relevant only in comparison of the experimental values and the perturbative calculations, and is irrelevant to the discussion on the existence of effective three- and four-body interactions.